The moon has a mass of 7.34 x 10²² kg and a radius of 1.74 x 10⁶ meters. If you have a mass of 66kg, how strong is the force between you and the moon?

1
Expert's answer
2022-01-05T14:30:08-0500
"F_g=\dfrac{GMm}{R^2},""F_g=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times7.34\times10^{22}\ kg\times66\ kg}{(1.74\times10^6\ m)^2},""F_g=106.7\ N."